# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v1

@@ B.4 Ablation Study on the Impact of grid size, $G$
(b) Experiments on MNIST and Fashion-MNIST.
@@
B.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
@@
To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. We use Algorithm 1 to project learned attention vectors to the probability simplex and compare Top-1 accuracies to the baseline model. We note that using $\ell_{1}$ projection does not increase the training time substantially. However, from Table 5, we observe that incorporating this algorithm in the Fourier-KArAt does not improve its performance; instead, the performance degrades significantly. Without applying Algorithm 1, the weights in each learnable unit can easily adjust and update accordingly.
B.6 Alternate Variants of Fourier KArAt
@@
B.6.1 Alternative Approaches to the Lower Rank Attention Structures
@@
In the main paper, we approximated the effect of the operator, ${\Phi^{i,j}}\in\mathbb{R}^{N\times N}$ casting it as the product of two rank-$r$ operators—one operator with learnable activation, $\widehat{\Phi},$ and the other is the learnable the linear transformation matrix, $W$ such that $\Phi=\widehat{\Phi}W$. Here, we abuse the notations for simplicity. However, a natural question is how many different configurations are possible with $\widehat{\Phi}$ and $W$ such that they can approximate the effect of $\Phi.$ Specifically, we use the following configurations: (i) [^2] This configuration was used in the main paper. $W\widehat{\Phi}(\mathcal{A}),$ where $\widehat{\Phi}\in\mathbb{R}^{r\times N}$ and $W\in\mathbb{R}^{N\times r}$, and (ii) $\widehat{\Phi}(W{\cal A}),$ where $W\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}\in\mathbb{R}^{N\times r}$. In the first configuration, $\widehat{\Phi}(\cdot)$ acts on each row of ${\cal A}$ to produce a $r$-dimensional vector, and then $W$ projects it back to an $N$-dimensional subspace. In the second configuration, $W$, first projects each row of ${\cal A}$ to produce a $r$-dimensional vector, and then $\widehat{\Phi}$ with learnable activation produces $N$-dimensional vectors.
@@
Primarily, we started with the full-rank operator $\Phi\in\mathbb{R}^{N\times N}$ and found their calculations are prohibitively expensive, regardless of the choice of the basis. Next, we conduct an ablation study to see which operator configuration works better; see Table 7. Our experiments show that the operator configuration (ii) demonstrates an inferior performance. We postulate that by down-projecting the $N$ dimensional attention row vector to a smaller dimension, $r$ loses adequate token-to-token interaction, and this fails to capture dependencies from significant attention units within a row. After that, from this limited information, learnable activation cannot significantly help the model’s performance.
B.6.2 Fourier KArAt and softmax—An Hybrid version
@@
We consider another variant where we mix the learnable activation and pre-defined activation in each head. E.g., the user can have 2 of the 3 attention heads in each encoder block in ViT-Tiny activated by Softmax and the third activated by the Fourier KArAt. With the idea of incorporating KAN as a way to replace softmax activation, we are curious to see if the ViT model performs better with a hybrid mode of activation; see Table 7 for results.
B.6.3 More Combinations
@@
We have also experimented with more configurations of Fourier KArAt in various permutations of the strategies mentioned above and have not found any significant combination in terms of performance. We also carefully design a particular training strategy where the attention operators ($\widehat{\Phi}$, $W$, and ${\Phi}$ wherever applicable) of Fourier KArAt are trained with a separate learning rate to alleviate the problem of the mismatch of the data passing through these learnable layers, as they consider each row of an attention matrix ${\cal A}$ as an input. This also attends to our observation of the models having gradient and loss explosion during training. However, we found this strategy to be of no improvement over the performance without the same.
B.7 B-Spline experiments
We test the performance of the B-spline KANs in the classification task on the CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky et al., 2009), MNIST (LeCun et al., 1998, 2010), and Fashion-MNIST (Xiao et al., 2017) datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 8 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works (Yu et al., 2024; Liu et al., 2025). Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range $[-I,I]$, layers, and width. Although the B-spline KANs yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets.

### v2

@@ B.3.1 Experiments with KANs Operated on B-Spline Basis
(b) Experiments on MNIST and Fashion-MNIST
@@
B.3.2 Ablation with the Hidden Dimension, $r$ in Fourier KArAt
While avoiding the computational overhead for computing $\Phi^{i,j}\in\mathbb{R}^{N\times N}$, we make use of the low-rank structure that the attention heads show (see Figure 6) by comparing different values of $r$. Particularly, we consider the values, $r=8,12,24$, on the Fourier KArAt variant of ViT-Base model, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$; see Table 16 for results on CIFAR-10. In this ablation, we observe that changing the hidden dimension has a negligible impact on the model’s performance. This can be explained by the sudden drop in the singular values, as shown in Figure 13. As long as $r$ remains greater than the sudden drop index, the model should not be impacted by the changing of $r$ except for changes in computational requirement; a higher $r$ would incur a higher compute time as the size of the operator $\widehat{\Phi}$ scales with $r$.
Model Acc.@1
CIFAR-10 CIFAR-100
ViT-Base 83.45 58.07
+ $G_{1}B$ 81.81 55.92
+ $G_{1}U$ 80.75 57.36
+ $G_{3}B$ 80.09 56.01
+ $G_{3}U$ 81.00 57.15
+ $G_{5}B$ 79.80 54.83
+ $G_{5}U$ 81.17 56.38
+ $G_{11}B$ 50.47 42.02
+ $G_{11}U$ 40.74 39.85
ViT-Small 81.08 53.47
+ $G_{1}B$ 79.00 53.07
+ $G_{1}U$ 66.18 53.86
+ $G_{3}B$ 79.78 54.11
+ $G_{3}U$ 79.52 53.86
+ $G_{5}B$ 78.64 53.42
+ $G_{5}U$ 78.75 54.21
+ $G_{11}B$ 77.39 52.62
+ $G_{11}U$ 78.57 53.35
ViT-Tiny 72.76 43.53
+ $G_{1}B$ 75.75 45.77
+ $G_{1}U$ 74.94 46.00
+ $G_{3}B$ 76.69 46.29
+ $G_{3}U$ 75.56 46.75
+ $G_{5}B$ 75.85
+ $G_{5}U$ 74.71
+ $G_{7}B$ 75.11
+ $G_{7}U$ 74.45
+ $G_{9}B$ 74.85
+ $G_{9}U$ 73.97
+ $G_{11}B$ 74.52
+ $G_{11}U$ 73.58
Model $r$ Acc.@1 on CIFAR-10
ViT-Base 58.07
+ $G_{3}B$ 24 80.54
+ $G_{3}U$ 24 80.81
+ $G_{5}B$ 24 77.99
+ $G_{5}U$ 24 80.52
+ $G_{3}B$ 12 80.09
+ $G_{3}U$ 12 81.00
+ $G_{5}B$ 12 79.80
+ $G_{5}U$ 12 81.17
+ $G_{3}B$ 8 80.76
+ $G_{3}U$ 8 80.40
+ $G_{5}B$ 8 79.79
+ $G_{5}U$ 8 80.83
Table 16: Comparing performance of Fourier KArAt for ViT-Tiny and ViT-Base with or without using $\ell_{1}$ projection in Algorithm 1. ViT-Tiny and ViT-Base use traditional softmax and do not require $\ell_{1}$ projection.
Model $\ell_{1}$ Projection Acc.@1
on CIFAR-10
ViT- Tiny 72.76
+ $G_{3}B$ 41.99
+ $G_{3}U$ 40.85
+ $G_{3}B$ 76.69
+ $G_{3}U$ 75.56
ViT-Base 83.45
+ $G_{3}B$ 47.44
+ $G_{3}U$ 46.11
+ $G_{3}B$ 80.09
+ $G_{3}U$ 81.00
Table 17: Complete compute requirement of ViT-Tiny+Fourier KArAt with extremely low hidden dimension training.
Model $r$ Acc.@1 Parameters GFLOPs GPU Memory
CIFAR-10 CIFAR-100
ViT-Tiny 72.76 43.53 5.53M 1.262 2.94GB
+ $G_{3}B$ 12 76.69 46.29 6.29M 1.425 7.48GB
4 75.90 45.76 5.80M 1.313 4.68GB
2 74.39 44.36 5.56M 1.285 4.01GB
+ $G_{3}U$ 12 75.56 46.75 5.59M 1.425 7.29GB
4 73.13 44.98 5.57M 1.313 5.08GB
2 71.56 42.43 5.55M 1.285 4.41GB
ViT-Base 83.45 58.07 85.81M 17.595 7.44GB
+ $G_{1}B$ 12 81.81 55.92 87.51M 17.847 17.36GB
4 81.20 55.29 86.41M 17.668 12.05GB
2 81.55 56.49 86.13M 17.623 9.48GB
+ $G_{1}U$ 12 80.75 57.36 85.95M 17.847 16.97GB
4 79.69 55.70 85.89M 17.668 12.11GB
2 80.36 54.04 85.87M 17.623 9.92GB
B.3.3 Enforcing an Extreme Low Rank Structure in Fourier KArAt
While our modular design of Fourier KArAt avoids the computational requirement of $\Phi^{i,j}$ by using a low hidden dimension $r$, it also enforces a low-rank structure in the attention matrix ${\cal A}^{i,j}$. In this context, we attempt to find the best possible value for $r$; see §B.3.2. However, we want to investigate how low the hidden dimension $r$ can be used without substantially compromising the performance of learnable Fourier KArAt. The significance of this question lies in the information bottleneck created by an extremely low hidden dimension $r$ that helps to understand the tradeoff between computing requirements and final trained model quality. To this end, we experiment with extreme low-rankness in the hidden dimension, $r=2,4$, and report our findings in Table 17. Although $r=2,4$ are not optimal, and we observe an insignificant drop in the performance across both datasets (CIFAR-10 and CIFAR-100), it comes at a reduced computing requirement, especially required VRAM; see GPU memory in Table 17. For instance, ViT-Tiny+$G_{3}B$ with $r=2$ outperforms the vanilla variant, only has a modest 0.03M parameters increment from the vanilla ViT-Tiny, albeit similar GFLOPs, but an extra 1.07GB of GPU memory usage. Compared to the $r=12$ variant, the memory usage is 3.47 GB lower, a total of 0.73M parameters less, but the relative decrease in accuracy is only 2.96% lower on the CIFAR-10 dataset. In ViT-Base backbones, the performance remains comparable, and even in the case of ViT-Base+$G_{1}B$ on CIFAR-100, the variant with $r=2$ surpasses the original variant with $r=12$. Although there is no concrete strategy, this observation supports our claim on a probable research direction involving parameter reduction (discussed in §6) to improve scalability in such over-parameterized models. This also indicates that for over-parameterized larger models, the scope of parameter reduction may find a model with learnable attention that can outperform its traditional counterpart across multiple datasets.
B.3.4 Ablation on the Impact of grid size, $G$ on Fourier KArAt
KANs are highly dependent on certain hyperparameters, and Fourier-KArAt has only one hyperparameter to tune the performance — grid size $G$. Thus, we perform extensive experiments involving grid size $G$ and present them in Table 16. We observe that each of the particular ViT models, in conjunction with particular Fourier-KArAt variants, has a typical $G$ value that brings out its best performance, and there is no universal value of $G$ to follow. When performing validation with ViT-Base+ variants on CIFAR-10 and CIFAR-100, the accuracy drops as the grid size passes a size after 5. However, this behavior is not persistent with the ViT-Small/Tiny+ variants; see Table 16333We did not perform gridsize 5, 7, 9, and 11 experiments with ViT-Tiny+KArAt as it was already outperforming the base ViT-Tiny with gridsize 1 and 3, and the experiments are extensively resource intensive..
Figure 11: Optimization path for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. We show the loss contours along with the trajectory of the optimizers. Perturbed contours indicate a corresponding non-smooth, spiky loss surface.
Figure 12: Vit-Tiny attention map characterization. Original image for inference (the center), the attention maps (top row), and contributing image regions (bottom row) for all three heads in ViT-Tiny: traditional MHSA (left) and Fourier KArAt $G_{3}B$ (right). The traditional MHSA sporadically focuses on fine-grained features of the primary object in the image. In contrast, the learnable attention in Fourier KArAt identifies the primary object features present significantly across all heads.
Table 18: Ablation study on the operator variants in Fourier KArAt. Note ${\widehat{\Phi}(W{\cal A})}$ and ${W\widehat{\Phi}({\cal A})}$ refer to order of using linear projector $W$ and operator $\widehat{\Phi}$, and $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ refers to two consecutive learnable operators $\widehat{\Phi}_{[1]}$ and $\widehat{\Phi}_{[2]}$.
Model Acc.@1 on CIFAR-10
$\widehat{\Phi}(W{\cal A})$ $W\widehat{\Phi}({\cal A})$ $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$
ViT-Base
+ $G_{11}B$ 41.02 50.47 30.14
+ $G_{11}U$ 47.74
ViT-Tiny
+ $G_{11}B$ 67.04 75.41 33.09
+ $G_{11}U$ 70.98 73.67 36.98
Model Low Rank Configurations No. of Distinct Heads Acc.@1 on
(for Fourier KArAt) softmax Fourier KArAt CIFAR-10
ViT-Base 12 0 83.45
ViT-Base + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 0 12 41.02
$\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 6 6 47.60
$\hat{\sigma}=W\widehat{\Phi}({\cal A})$ 0 12 50.47
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 0 12 30.14
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 6 6 30.24
ViT-Tiny 3 0 72.76
ViT-Tiny + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}_{[1]}(\widehat{\Phi}_{[2]}({\cal A})$ 0 3 33.09
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 1 2 32.51
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 2 1 34.73
ViT-Tiny + $G_{11}U$ $\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 0 3 36.98
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 1 2 33.12
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 2 1 70.57
B.3.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
@@
To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project it onto the $\ell_{1}$-unit ball. We use Algorithm 1 to project learned attention vectors to the probability simplex and compare Top-1 accuracies to the baseline model. We note that using $\ell_{1}$ projection does not substantially increase the training time. However, from Table 16, we observe that incorporating this algorithm in the Fourier-KArAt does not improve its performance; instead, the performance significantly degrades. Algorithm 1 can be used with the choice of any basis function. However, we cannot comment on the role of $\ell_{1}$-projection in KArAt’s performance with basis functions other than Fourier.
B.3.6 More Discussion on Blockwise and Universal Configuration
Although the primary inspiration behind the universal configuration comes from the shared basis functions in KAT [80], it aligns with the fact that softmax is a fixed function conventionally being used across all heads of all encoder layers in the ViT architectures. This raises the question of whether the token-to-token interaction in the self-attention can be modeled using a single universal function, or each instance of attention operations needs dynamic modeling using learnable functions. To investigate further, we compared the learned coefficients $\{a_{pqm}$, $b_{pqm}\}$ from Equation (1) and the weight matrix $W$ of the linear projector following the learnable basis, for each head across all the encoder layers in ViT-Tiny+Fourier KArAt. Thus, we obtain 36 sets (3 heads for each of the 12 layers) of such coefficients and weights from the model with blockwise configuration $G_{3}B$, and 3 sets from the model with universal configuration $G_{3}U$. Although mostly centered at zero, the distribution of each of these sets of parameters varies significantly from the others. The difference is mostly visible in the presence of long tails and the length of the tails. This characteristic is present not only in the blockwise configuration, but the three heads of the universal configuration also show a similar pattern.
(a) Before attention activation
(b) After attention activation
Figure 13: Spectral decomposition of the attention matrix for ViT-Tiny on CIFAR-10 dataset with traditional softmax attention and our learnable Fourier KArAt. The traditional softmax attention and our learnable Fourier KArAt have almost similar low-rank structure, before activation functions are used.
B.4 Alternate Variants of Fourier KArAt
@@
B.4.1 Alternative Approaches to the Lower Rank Attention Structures
@@
In the main paper, we approximated the effect of the operator, ${\Phi^{i,j}}\in\mathbb{R}^{N\times N}$ casting it as the product of two rank-$r$ operators—one operator with learnable activation, $\widehat{\Phi},$ and the other is the learnable the linear transformation matrix, $W$ such that $\Phi=\widehat{\Phi}W$. Here, we abuse the notations for simplicity. However, a natural question is how many different configurations are possible with $\widehat{\Phi}$ and $W$ such that they can approximate the effect of $\Phi.$ Specifically, we use the following configurations — (i) [^4] This configuration was used in the main paper. $W\widehat{\Phi}(\mathcal{A}),$ where $\widehat{\Phi}\in\mathbb{R}^{r\times N}$ and $W\in\mathbb{R}^{N\times r}$, (ii) $\widehat{\Phi}(W{\cal A}),$ where $W\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}\in\mathbb{R}^{N\times r}$, and (iii) $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$, where $\widehat{\Phi}_{[1]}\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}_{[2]}\in\mathbb{R}^{N\times r}$ are low-rank operators for learning activation functions. In the first configuration, $\widehat{\Phi}(\cdot)$ acts on each row of ${\cal A}$ to produce an $r$-dimensional vector, and then $W$ projects it back to an $N$-dimensional subspace. In the second configuration, $W$, first projects each row of ${\cal A}$ to produce an $r$-dimensional vector, and then $\widehat{\Phi}$ with learnable activation produces $N$-dimensional vectors. In the third configuration, $\widehat{\Phi}_{[1]}$ first produces $r$-dimensional vector from rows of ${\cal A}$ by learning activations, and $\widehat{\Phi}_{[2]}$ obtains $N$-dimensional vectors by learning a second level of activations.
@@
Primarily, we started with the full-rank operator $\Phi\in\mathbb{R}^{N\times N}$ and found its computations to be prohibitively expensive, regardless of the choice of the basis. Next, we conduct an ablation study to see which operator configuration works better; see Table 19. Our experiments show that the operator configuration (ii) demonstrates an inferior performance. We postulate that by down-projecting the $N$-dimensional attention row vector to a smaller dimension, $r$ loses adequate token-to-token interaction, and this fails to capture dependencies from significant attention units within a row. After that, from this limited information, learnable activation cannot significantly help the model’s performance. On the other hand, (iii) also fails to perform due to its inability to model the attention well, despite having refined information, from the intermediate $r$-dimensional subspace. Overall, apart from configuration (i), the performances of the other configurations were inconsistent over multiple experiments. Also, they lag in training stability, particularly configuration (iii). Considering these observations, we proceeded to find the best possible grid size $G$ only with the configuration (i).
B.4.2 Fourier KArAt + softmax Attention —A Hybrid version
@@
We consider another variant where we mix the learnable and pre-defined activation in each head. E.g., it is possible to have 2 of the 3 attention heads in each encoder block in ViT-Tiny activated by softmax and the third activated by the Fourier KArAt. With the idea of incorporating KAN to replace softmax activation, we are curious to see if the ViT model performs better with a hybrid mode of activation; see Table 19 for results.
B.4.3 More Combinations
@@
We have also experimented with more configurations of Fourier KArAt in various permutations of the strategies mentioned above and have not found any significant combination in terms of performance. We also carefully design a particular training strategy where the attention operators ($\widehat{\Phi}$, $W$, and ${\Phi}$ wherever applicable) of Fourier KArAt are trained with a separate learning rate to alleviate the problem of the mismatch of the data passing through these learnable layers, as they consider each row of an attention matrix ${\cal A}$ as an input. The models have a gradient and loss explosion during training.
B.5 Spectral Analysis of Attention
Although we cannot comment on the generalizability by studying the distribution of the weights, from Table 2, we realize all the KArAt variants have more parameters than their vanilla counterparts. Therefore, it would be interesting to see how their attention matrices behave. As discussed in §4, the attention matrices in traditional MHSA have a low-rank structure. Following that study, we verified the apparent low-rank structure that Fourier KArAt’s learned attention matrices possess.
We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) for any attention matrix of shape $197\times 197$, where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $[\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads.
To investigate the inherent low-rank structure of attention, we plot the natural logarithm of singular values of attention matrices, $\sigma_{i}$, before and after attention activation in Figure 13. From Figure 13, we observe that the traditional softmax attention and our learnable Fourier KArAt have almost similar low-rank structures. However, Figure 13 shows that the traditional MHSA has significantly larger singular values than its KArAt variant, $G_{3}B$. It can also be noticed that before the activation, both traditional MHSA and Fourier KArAt feature a sharp drop in singular values between the 50th and 75th indices. This sharp drop in singular values vanishes in softmax attention, indicating a normalization. However, due to the hidden dimension $r$, Fourier KArAt enforces a much lower rank than the traditional MHSA.
Additionally, we observe from the weight distributions in Figure 3 that Fourier KArAt variants have more entries close to zero than their ViT counterparts. This, along with the low-rankness, can inspire low-rank + sparse training strategies of Fourier KArAT.